OpenStudy (anonymous):

Solve x2 + 6x + 7 = 0. can anyone help

OpenStudy (anonymous):

\[\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\] can you insert for variables a=1 b=6 c=7

OpenStudy (anonymous):

so the answer would be

OpenStudy (anonymous):

close you're missing a number though

OpenStudy (anonymous):

would it be over 2

OpenStudy (anonymous):

\[\frac{ -6 \pm \sqrt {6^2-4(1)(7)} }{ 2(1) }\] \[\frac{ -6 \pm \sqrt {36-28} }{ 2 }\] \[\frac{ -6 \pm \sqrt {8} }{ 2 }\]

OpenStudy (anonymous):

What does 8 factor out to?

OpenStudy (anonymous):

the result product would be 2